OpenStudy (anonymous):

Create a polynomial with theese qualities: Degree 4, 2 positive real zeros, 0 negative real zeros, 2 complex zeros.

OpenStudy (campbell_st):

its simply P(x) = (x - 3)(x -2)(x -2i)(x + 2i) real positive zeros 3, 2 complex zeros, -2i, 2i

OpenStudy (campbell_st):

you just need to expand it out

OpenStudy (anonymous):

Thanks, but how exactly did you figure that out??

OpenStudy (campbell_st):

well start by distributing (x -3)(x -2) = and (x -2i)(x+2i) = which is the difference of 2 squares then multiply the 2 answers

OpenStudy (campbell_st):

well its easy to figure out. I just chose some numbers roots of polynomials are the values that make it zero P(x) = (x + a)(x -b) the roots are -a and b so if you are asked for real positive roots then it would be x = a, b so the binomial factors are P(x) = (x -a)(x-b) complex roots involve i because i^2 = -1 as an example if P(x) = x^2 + 9 to find the zeros you have x^2 + 9 = 0 solving for x you get x^2 = -9 so using -1 = i^2 x^2 = 9i^2 now you can find roots at -3i and 3i reallY i just chose random numbers that met the conditions.